George is 5 times as old as his son. In 15 years he will be only twice as old as his son. How old will his son be in 15 years?
Answer by stanbon(75887)

George is 5 times as old as his son. In 15 years he will be only twice as old as his son. How old will his son be in 15 years?
---------------------------------------------
Equations:
g = 5s
(g+15) = 2(s+15)
-----------------------------
Substitute for "g" and solve for "s":
5s+15 = 2s + 30
3s = 15
s = 5 (son's age NOW)
s+15 = 20 (san's age in 15 years)
===================================
Cheers,
Stan H.